Conditions (18h30 MST) : - clear but clouds coming in from the West

- humidity 50%

- seeing 1"-1.2"

18:50 icmslock issue

19:07 sent the preset on acq star again

19:25 clouds... std spectra has 4x less counts than wednesday

19:26 downtime: somebody needs to shut down a valve on the roof

19:32 moving to CSWA164

19:47 started gz obs

20:00 moving to CSWA116 seeing 1.5"

clouds coming in from the west, last exposures are probably not usable

20:26 moving to CSWA139 seeing 1.6"

20:32 started exposure but some cirrus and seeing 1.8"...

20:45 difficulty in guiding because of clouds, switching to backup QSO J074154

20:56 clouds : guiding lost and acquisition star not seen

21:26 overcast

22:06 mirror panicked, hard to collimate given present conditions (clouds)

23:58 small hole in the clouds, we are moving to CSWA139, seeing is ~1.2"

00:39 now on CSWA16 seeing 1.1"

01:00 moving to CSWA141

done

01:29 moving to A1703

01:48 started science 3*1200s, got some clouds during acquisition, seeing 1.2-1.3"

02:54 moved to CSWA13

data not good, clouds

03:04 overcast, humidity is rising, we are now closed, going to zenith to take flats

05:28 humidity 100%, some snowflakes, MODS has been put to bed

-- BenjaminClement - 27 Jan 2014
Topic revision: r4 - 27 Jan 2014, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback